# The opaque law of artificial intelligence

### v1

@@ 1 The black box poblem
Let’s make a brief example: a cat is a creature with free will, such as humans, and by so I mean that they both are indeterministic objects, in the sense that their actions are not completely determined by the inputs that are given to them. If a cat broke a flower vase, it is an action of its own and under a legal point of view the only possible responsible can be the owner. That’s because the owner of a cat has the obligation of stay in control of his pet, for the reason that our legal systems are designed on human values which are not relevant for animal’s action. Animal, in fact, are object of law, not subject. Even AI algorithms are object of law, as stated by the AI act, but their position is completely different under a substantial point of view: AI can’t act independently, in opposite of a living creature of some degree of complexity. For example you can expect by a natural language processing (NLP) model [^3] NLP are “Techniques used by large language models to understand and generate human language, including text classification and sentiment analysishttps://www.nytimes.com/article/ai-artificial-intelligence-glossary.html. a lot of different and strange words as output if given the right prompt, but at the same time that AI will not do other things except give you back an answer, because it is not independent and only respond at input given by humans.
@@ 2 Generative AI
Natural language processing is a flourishing field in the process of AI development. NLP is a complex activity which does not involve only computer science but it is a major component of law, politics and philosophy; it could be said that natural language processing is the architecture of humanistic culture and, for decades, it was a challenge for machines to understand and compute human natural language for its informality e semantic complexity. The existence of formal rules that may be written by computers algorithm, by which they became capable of formulating natural language, is the accomplishment of at least seventy years of research, since the publishing of Computing machinery and intelligence[29] where Alan Turing opened for the first time in history to the possibility that a machine could be capable of mimicking natural human language.
@@ 3 An interpretation of the halting problem
Let’s start to consider an argument from computablity theory: the halting problem. The halting problem or decision problem (Entscheidunsproblem) is the problem which both Turing and Church approached and which led to the development of the concepts of computation and effective calculability, which are substantially equivalent. Originally the Entscheidunsproblem was about the possibility of finding a general method to verify if a logical expression is true or not, with a yes/no final ending[3]; Turing and Church, with their theses, demonstrated that it is impossible to create a universal computational method which can decide for any given logical expression if it is true, but they found two different methods, the lambda calculus and the Turing machine, to decide whether for a specific statement there is an absolute mechanical procedure, namely an algorithm, to state its satisfiability [^28] With the development of computational complexity theory, the problem has been refined. If a fragment of first-order logic is decidable for satisfiability, then indeed there is an absolutely mechanical procedure, that is an algorithm, for deciding the satisfiability or unsatisfiability of any given sentence” from pag. 7 of Börger, Egon, Erich Grädel, and Yuri Gurevich. The classical decision problem. Springer Science & Business Media, 2001..
After the Turing-Church thesis, the Entscheidunsproblem has started to be presented as the halting problem, probably influenced by the major evidence that the Turing machine had on the topic. Instead of searching for a general method to verify the satisfiability of every logical expression, researchers started to search for a single method to verify the satisfiability of a single logical expression[3].

### v2

@@ 1 The black box problem
Let’s make a brief example: a cat is a creature with free will, such as humans, and by so I mean that they both are nondeterministic objects, in the sense that their actions are not completely determined by the inputs that are given to them. If a cat broke a flower vase, it is an action of its own and under a legal point of view the only possible responsible can be the owner. That’s because the owner of a cat has the obligation of stay in control of his pet, for the reason that our legal systems are designed on human values which are not relevant for animal’s action. Animal, in fact, are object of law, not subject. Even AI algorithms are object of law, as stated by the AI act, but their position is completely different under a substantial point of view: AI can’t act independently, in opposite of a living creature of some degree of complexity. For example you can expect by a natural language processing (NLP) model [^3] NLP are “Techniques used by large language models to understand and generate human language, including text classification and sentiment analysishttps://www.nytimes.com/article/ai-artificial-intelligence-glossary.html. a lot of different and strange words as output if given the right prompt, but at the same time that AI will not do other things except give you back an answer, because it is not independent and only respond at input given by humans.
@@ 2 Generative AI
Natural language processing is a flourishing field in the process of AI development. NLP is a complex activity which does not involve only computer science but it is a major component of law, politics and philosophy; it could be said that natural language processing is the architecture of humanistic culture and, for decades, it was a challenge for machines to understand and compute human natural language for its informality and semantic complexity. The existence of formal rules that may be written by computers algorithm, by which they became capable of formulating natural language, is the accomplishment of at least seventy years of research, since the publishing of Computing machinery and intelligence[29] where Alan Turing opened for the first time in history to the possibility that a machine could be capable of mimicking natural human language.
@@ 3 An interpretation of the halting problem
Let’s start to consider an argument from computablity theory: the halting problem. The halting problem or decision problem (Entscheidungsproblem) is the problem which both Turing and Church approached and which led to the development of the concepts of computation and effective calculability, which are substantially equivalent. Originally the Entscheidungsproblem was about the possibility of finding a general method to verify if a logical expression is true or not, with a yes/no final ending[3]; Turing and Church, with their theses, demonstrated that it is impossible to create a universal computational method which can decide for any given logical expression if it is true, but they found two different methods, the lambda calculus and the Turing machine, to decide whether for a specific statement there is an absolute mechanical procedure, namely an algorithm, to state its satisfiability [^28] With the development of computational complexity theory, the problem has been refined. If a fragment of first-order logic is decidable for satisfiability, then indeed there is an absolutely mechanical procedure, that is an algorithm, for deciding the satisfiability or unsatisfiability of any given sentence” from pag. 7 of Börger, Egon, Erich Grädel, and Yuri Gurevich. The classical decision problem. Springer Science & Business Media, 2001..
After the Turing-Church thesis, the Entscheidungsproblem has started to be presented as the halting problem, probably influenced by the major evidence that the Turing machine had on the topic. Instead of searching for a general method to verify the satisfiability of every logical expression, researchers started to search for a single method to verify the satisfiability of a single logical expression[3].